OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66. 70 72 111 114

OpenStudy (anonymous):

Is this urgent?

OpenStudy (anonymous):

kind of yeah.

OpenStudy (anonymous):

the quarter is ending soon and i have a few assignments that have a couple problems i just cant get...

OpenStudy (anonymous):

Okay speed round. So there are 27-1=26 differences between a1 and a27. The difference between a1 and a27 is 66-(-12)=78. that means the n (the difference between each term in the sequence is 78/26=3. Can you take it from here?

OpenStudy (anonymous):

I will be back...

OpenStudy (anonymous):

where do you get the -12?

OpenStudy (anonymous):

Isn't that a1?